OpenStudy (anonymous):

find the zeros using factoring for y=x^2+x-12

OpenStudy (lalaly):

x^2+x-12=0 (x+4)(x-3)=0 x=-4 , x=3

OpenStudy (anonymous):

Ummm , thats not one of the choices

OpenStudy (lalaly):

what are the choices

OpenStudy (akshay_budhkar):

x=-4 or 3 is right....else your question is wrong taylerr

OpenStudy (anonymous):

Nvm i was looking at the wrong question but thankyou both